A friend asks you how much pressure is in your car tires. Unfortunately, you have misplaced your gauge but confident in your knowledge of physics you tell your friend “I can figure it out”. From the owner’s manual you find that the car’s mass is 1500 kg and that the tire manufacturer recommends 31 psi. Assuming that each tire supports one-fourth of the weight of the car, you use a ruler to find that the tires are 15 cm wide and the flattened segment of the tire in contact with the road is 15 cm long. What answer do you give your friend? Express your answers in pounds per square inch (psi)
Answer #1

Pressure, P = F/A

For each tire, F= mg = (1500/4) x 9.8 = 3675 N

Area under pressure A = 0.15 x 0.15 = 0.0225 m2

Hence P = 3675/0.0225 = 163333.33 Pa

We have 1Pa = 0.000145psi

Hence P = 163333.33 x 0.000145

Answer:- P = 23.68 psi

Hope the answer helps you. All the best for your studies ?.